# Parameter estimation for Gibbs distributions

### v9

@@ 8 Lower bounds on sample complexity
In this section, we show lower bounds for the sample complexity needed to solve ${P^{{\varepsilon}}_{\tt ratio-pt}}$ and ${P^{\delta,\varepsilon}_{\tt count}}$ in the black-box model. The specification of problem ${P^{\delta,\varepsilon}_{\tt count}}$, while useful from an algorithmic point of view, has some restrictive choices for its normalization and error bounds; to get more general lower bounds for count estimation, we consider a relaxed problem called ${P^{\delta,\varepsilon}_{\tt count-weak}}$, namely, to compute a vector $\hat{c}\in\mathbb{R}_{\geq 0}^{\mathcal{F}}$ satisfying the following property:
$|\log\frac{\hat{c}_{x}}{\hat{c}_{y}}-\log\frac{c_{x}}{c_{y}}|\leq\varepsilon\qquad\qquad\text{ for all pairs $x,y$ with $\Delta(x),\Delta(y)\geq\delta$}$
@@
Our strategy for the lower bounds is to construct a target instance $c$ surrounded by an envelope of alternate instances, such that solving ${P^{{\varepsilon}}_{\tt ratio-pt}}$ or ${P^{\delta,\varepsilon}_{\tt count-weak}}$ on an unknown instance $\tilde{c}$ allows us to distinguish between whether $\tilde{c}$ is an alternate instance or the base instance. Specifically, let us choose $d$ vectors $\lambda^{(1)},\dots,\lambda^{(d)}\in\mathbb{R}^{\mathcal{F}}$. We will define $2d$ alternate instances $r=-d,\dots,-1,1,\dots,d$ which straddle the base instance $c=c^{(0)}$, defined as:
[…]
For $r=-d,\dots,d$, we define $\mu_{\beta}^{(r)}$ to be the Gibbs distribution for instance $c^{(r)}$, and $Z^{(r)}$ to be its partition function, and $z^{(r)}=\log Z^{(r)}$, etc. When the superscript is omitted, we always refer to the base instance $c=c^{(0)}$, e.g. we write $z(\beta)=z^{(0)}(\beta),\Delta(x)=\max_{\beta\in[\beta_{\min},\beta_{\max}]}\mu_{\beta}^{(0)}(x)$.
@@
Let $\mathfrak{A}$ be an algorithm which generates queries $\beta_{1},\dots,\beta_{T}\in[\beta_{\min},\beta_{\max}]$ and receives values $x_{1},\dots,x_{T}$, where each $x_{i}$ is drawn from $\mu_{\beta_{i}}$. At some point the procedure stops and outputs TRUE or FALSE. The queries $\beta_{i}$ and the stopping time $T$ may be adaptive and may be randomized.
Let $\gamma\in(0,1/4)$. If $\mathfrak{A}$ outputs FALSE on input $c^{(0)}$ with probability at least $1-\gamma$ and outputs FALSE on inputs $c^{(-d)},\dots c^{(-1)},c^{(1)},\dots,c^{(d)}$ with probability at least $1-\gamma$, then $\mathfrak{A}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})$ when run on instance $c^{(0)}$.
@@
Any finite run of $\mathfrak{A}$ can be described by a random vector $X=((\beta_{1},x_{1}),\ldots,(\beta_{T},x_{T}),\sigma)$, where $\sigma\in\{\text{TRUE},\text{FALSE}\}$. The probability measure ${\mathbb{P}}^{(i)}(\cdot)$ over such runs $X$ on an instance $c^{(r)}$ can be decomposed as:
@@
Let $\mathcal{X}$ denote the set of runs of length at most $\tau=\frac{2d\log(1/(2\gamma))}{\Psi}$. Partition $\mathcal{X}$ into sets of runs $\mathcal{X}^{\text{T}},\mathcal{X}^{\text{F}}$ which output TRUE and FALSE respectively. For any run $X\in\mathcal{X}$ of length $T$, Eq. (24) gives:
$\frac{1}{2d}\sum_{r\in R}\psi^{(r)}(X)\geq\Bigl{(}\prod_{r\in R}\psi^{(r)}(X)\Bigr{)}^{1/(2d)}\geq\psi^{(0)}(X)e^{-T\Psi/2d}\geq 2\gamma\psi^{(0)}(X)$ (25)
where the first bound comes from the inequality between arithmetic and geometric means, and the last inequality holds since $T\leq\tau$.
By hypothesis of the lemma, we have ${\mathbb{P}}^{(r)}(\mathcal{X}^{\text{T}})<\gamma$ for all $r\in R$. Using Eq. (25), this implies
${\mathbb{P}}^{(0)}(\mathcal{X}^{\text{T}})=\negthinspace\int_{{\cal X}^{\text{T}}}\negthinspace\psi^{(0)}(X)\>\operatorname{d}\!{\mu}^{\mathfrak{A}}X\leq\frac{1}{4d\gamma}\sum_{r\in R}\int_{{\cal X}^{\text{T}}}\negthinspace\psi^{(r)}(X)\>\operatorname{d}\!{\mu}^{\mathfrak{A}}X=\frac{1}{4d\gamma}\sum_{r\in R}{\mathbb{P}}^{(r)}(\mathcal{X}^{\text{T}})\leq\frac{1}{4d\gamma}\sum_{r\in R}\gamma=\frac{1}{2}.$
By hypothesis, we have ${\mathbb{P}}^{(0)}(\mathcal{X}^{\text{F}})<\gamma$. So ${\mathbb{P}}^{(0)}(T>\tau)=1-{\mathbb{P}}^{(0)}(\mathcal{X}^{\text{T}})-{\mathbb{P}}^{(0)}(\mathcal{X}^{\text{F}})\geq 1-\gamma-\frac{1}{2}\geq\frac{1}{4}$. The expected cost of $\mathfrak{A}$ on $c^{(0)}$ is at least $\frac{1}{4}\cdot\tau=\Omega(\frac{d\log(1/\gamma)}{\Psi})$. ∎
@@
Suppose that $z^{(0)}(\beta_{\min},\beta_{\max})\leq q-2\kappa$.
(a) If $|z(\beta_{\min},\beta_{\max})-z^{(r)}(\beta_{\min},\beta_{\max})|>2\varepsilon$ for all $r=\pm 1,\dots,\pm d$, then any algorithm for ${P^{{\varepsilon}}_{\tt ratio-pt}}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})$ on $c$.
(b) If for each $r=1,\dots,d$ there are $x,y\in\mathcal{F}$ with $\Delta(x),\Delta(y)\geq\delta e^{2\kappa}$ and $|\lambda^{(r)}_{x}-\lambda^{(r)}_{y}|>2\varepsilon$, then any algorithm for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})$ on $c$. (We refer to $x,y$ as the witnesses for index $r$).
@@
As $|\log c^{(r)}_{x}-\log c_{x}|\leq\kappa$ for all $x$, we have $\frac{Z^{(r)}(\beta)}{Z(\beta)}\in[e^{-\kappa},e^{\kappa}]$ and $\mu^{(r)}_{\beta}(x)=\frac{c_{x}^{(r)}e^{\beta x}}{Z^{(r)}(\beta)}\geq\mu_{\beta}(x)e^{-2\kappa}$. So $z^{(r)}(\beta_{\min},\beta_{\max})\leq q$ for all $r=\pm 1,\dots,\pm d$. In particular, algorithms for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ or ${P^{{\varepsilon}}_{\tt ratio-pt}}$, which take as input the upper-bound value $q$, must also work for all instances $c^{(-d)},\dots,c^{(+d)}$. Furthermore, such algorithms can be used to distinguish $c^{(0)}$ from $c^{(\pm 1)},\dots,c^{(\pm d)}$. Namely, given a solution $\hat{Q}(\beta_{\max})$ for ${P^{{\varepsilon}}_{\tt ratio-pt}}$, we output TRUE if and only if $|\log\hat{Q}(\beta_{\max})-\log Q^{(0)}(\beta_{\max})|\leq\varepsilon$. Likewise, given a solution $\hat{c}$ to ${P^{\delta,\varepsilon}_{\tt count-weak}}$, we output TRUE if and only if for all $r=1,\dots,d$ the witnesses $x,y$ satisfy $|\log\frac{\hat{c}_{x}}{\hat{c}_{y}}-\log\frac{c_{x}}{c_{y}}|\leq\varepsilon$. Note that $\Delta^{(\pm r)}(x),\Delta^{(\pm r)}(y)\geq\delta$ and so $x,y$ can serve as witnesses for both instances $c^{(+r)}$ and $c^{(-r)}$. ∎
By applying Corollary 46 to carefully constructed instances, we will show the following results:
@@
Let $n\geq n_{0},q\geq q_{0},\varepsilon<\varepsilon_{0},\delta<\delta_{0},\gamma<1/4$ for certain absolute constants $n_{0},q_{0},\varepsilon_{0},\delta_{0}$. There are problem instances $\mu^{(0)},\dots\mu^{(d)}$ which satisfy the given bounds $n$ and $q$ such that:
@@
(c) ${P^{{\varepsilon}}_{\tt ratio-pt}}$ requires cost $\Omega(\frac{\min\{q,n^{2}\}}{\varepsilon^{2}}\log\frac{1}{\gamma})$, and $\mu$ is log-concave.
@@
(e) ${P^{{\varepsilon}}_{\tt ratio-pt}}$ requires cost $\Omega(\frac{q}{\varepsilon^{2}}\log\frac{1}{\gamma})$.
Theorem 47 immediately implies Theorems 2 and 4.
The lower bound in terms of $\delta$ for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ is rather trivial, and does not really depend on properties of Gibbs distributions. We summarize it in the following result:
@@
We next turn to the non-trivial lower bounds in terms of $n$ and $q$.
8.1 Bounds in the log-concave setting
For this construction, take $d=1,\beta_{\min}=0,\beta_{\max}=n$, and define $c_{0},\dots,c_{n}$ to be the coefficients of the polynomial $g(x)=\prod_{k=0}^{n-1}(e^{k}+x)$, and set $\lambda^{(1)}_{k}=k\nu$ for $k=0,\dots,n$ where $\nu=10\varepsilon/n$. Since $g(x)$ is a real-rooted polynomial, and counts $c^{(\pm 1)}$ are derived by an exponential shift, the counts $c^{(0)},c^{(+1)},c^{(-1)}$ are log-concave [5].
Here $\prod_{k=0}^{n-1}(e^{k}+x)=c_{0}+c_{1}x+\dots+c_{n}x^{n}$, and $c_{0}=\prod_{k=0}^{n-1}e^{k}$ and $c_{n}=1$, and we have
$Z(\beta)=\sum_{x}c_{x}e^{\beta x}=g(e^{\beta x})=\prod_{k=0}^{n-1}(e^{k}+e^{\beta}),\quad\text{and}\quad Z^{(\pm 1)}(\beta)=\sum_{x}c_{x}e^{\pm\nu}e^{\beta x}=Z(\beta\pm\nu)$
For intuition, consider independent random variables $X_{0},\dots,X_{n-1}$, wherein each $X_{i}$ is Bernoulli-$p_{i}$ for $p_{i}=\frac{e^{\beta}}{e^{i}+e^{\beta}}$. Then $\mu_{\beta}$ is the probability distribution on random variable $X=X_{0}+\dots+X_{n-1}$. Distributions $\mu^{(\pm 1)}$ are derived by slightly shifting the probabilities $p_{i}$.
@@ 8.1 Bounds in the log-concave setting
$\mu_{0}(0)$ $=\frac{c_{0}}{Z(0)}=\prod_{k=0}^{n-1}\frac{e^{k}}{e^{k}+1}\geq\prod_{k=0}^{n-1}\exp(-e^{-k})\geq\exp(-\sum_{k=0}^{\infty}e^{-k})=e^{-e/(e-1)}\geq 1/5.$
@@
For $n\geq n_{0},q>q_{0},\varepsilon<1/100$, where $n_{0},q_{0}$ are absolute constants, the instance can be constructed so that both ${P^{{\varepsilon}}_{\tt ratio-pt}}$ or ${P^{1/10,\varepsilon}_{\tt count-weak}}$ require cost $\Omega(\frac{\min\{q,n^{2}\}\log\frac{1}{\gamma}}{\varepsilon^{2}})$.
@@
By reducing $n$ as needed, and taking $n_{0},q_{0}$ sufficiently large, it suffices to show this under the assumption $10100<n<\sqrt{q}/$ 10. We use Corollary 46 for both problems. Here $\kappa=n\nu=10\varepsilon$, and so $z(\beta_{\min},\beta_{\max})+2\kappa\leq n^{2}+1/5\leq q$ for $q_{0}$ sufficiently large.
For ${P^{{\varepsilon}}_{\tt ratio-pt}}$ we have $|z(\beta_{\min},\beta_{\max})-z^{(\pm 1)}(\beta_{\min},\beta_{\max})|>2\varepsilon$ as required.
For ${P^{1/10,\varepsilon}_{\tt count-weak}}$ take witnesses $x=0,y=n$. We have $\Delta(x),\Delta(y)\geq 1/5\geq 1/10\cdot e^{2\kappa}$ and we have $|\lambda^{(1)}_{x}-\lambda^{(1)}_{y}|=n\nu=10\varepsilon$ as required.
So both problems require cost $\Omega(\frac{\log(1/\gamma)}{\Psi})=\Omega(\frac{\log(1/\gamma)n^{2}}{\varepsilon^{2}})$. ∎

### v11

@@ 8 Lower bounds on sample complexity
In this section, we show lower bounds for the sample complexity of ${P^{{\varepsilon}}_{\tt ratio-point}}$ and ${P^{\delta,\varepsilon}_{\tt count}}$ in the black-box model. At this point, we note that the specification of problem ${P^{\delta,\varepsilon}_{\tt count}}$, while useful from an algorithmic point of view, has some restrictive choices for its normalization and error terms. To get lower bounds which do not depend on these details, we consider a relaxed count-estimation problem called ${P^{\delta,\varepsilon}_{\tt count-weak}}$, namely, to compute a vector $\hat{c}\in\mathbb{R}_{\geq 0}^{\mathcal{F}}$ satisfying the following property:
$\Bigl{|}\log\frac{\hat{c}_{x}}{\hat{c}_{y}}-\log\frac{c_{x}}{c_{y}}\Bigr{|}\leq\varepsilon\qquad\qquad\text{ for all pairs $x,y$ with $\Delta(x),\Delta(y)\geq\delta$}$
@@
Our strategy for the lower bounds is to construct a target instance $c=c^{(0)}$ surrounded by an envelope of alternate instances which can be distinguished by solving ${P^{{\varepsilon}}_{\tt ratio-point}}$ or ${P^{\delta,\varepsilon}_{\tt count-weak}}$. Specifically, for vectors $\lambda^{(1)},\dots,\lambda^{(d)}\in\mathbb{R}^{\mathcal{F}}$, we will define $2d$ alternate instances $c^{(+1)},c^{(-1)},\dots,c^{(+d)},c^{(-d)}$ by:
$c^{(+r)}_{x}=e^{\lambda^{(r)}_{x}}c_{x},\qquad c^{(-r)}_{x}=e^{-\lambda^{(r)}_{x}}c_{x}\qquad\qquad\text{for $r=1,\dots,d$ and $x\in\mathcal{F}$}$
Thus, these $2d$ instances are arranged in $d$ pairs which “straddle” the base instance $c^{(0)}$.
For each instance $c^{(r)}$, we define $\mu_{\beta}^{(r)}$ to be the Gibbs distribution, and $Z^{(r)}$ to be its partition function, and $z^{(r)}=\log Z^{(r)}$, etc. When the superscript is omitted, we always refer to the base instance $c=c^{(0)}$, e.g. we write $z(\beta)=z^{(0)}(\beta),\Delta(x)=\max_{\beta\in[\beta_{\min},\beta_{\max}]}\mu_{\beta}^{(0)}(x)$.
@@
Let $\mathfrak{A}$ be an algorithm which generates queries $\beta_{1},\dots,\beta_{T}\in[\beta_{\min},\beta_{\max}]$ and receives values $x_{1},\dots,x_{T}$, where each $x_{i}$ is drawn from $\mu_{\beta_{i}}$. At some point the procedure stops and outputs YES or NO. The queries $\beta_{i}$ and the stopping time $T$ may be adaptive and may be randomized.
If $\mathfrak{A}$ outputs YES on instance $c^{(0)}$ with probability at least $1-\gamma$ and outputs NO on each instance $c^{(-d)},\dots c^{(-1)},c^{(1)},\dots,c^{(d)}$ with probability at least $1-\gamma$, then $\mathfrak{A}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})$ when run on instance $c^{(0)}$.
@@
Any finite run of $\mathfrak{A}$ can be described by a random vector $X=((\beta_{1},x_{1}),\ldots,(\beta_{T},x_{T}),\sigma)$, where $\sigma\in\{\text{YES},\text{NO}\}$. The probability measures ${\mathbb{P}}^{(r)}(\cdot)$ over such runs on $c^{(r)}$ can be decomposed as:
@@
Let $\mathcal{X}$ denote the set of runs of length at most $\tau=\frac{2d\log(1/(4\gamma))}{\Psi}$. For any run $X\in\mathcal{X}$, Eq. (24) gives:
$\frac{1}{2d}\sum_{r\in R}\psi^{(r)}(X)\geq\Bigl{(}\prod_{r\in R}\psi^{(r)}(X)\Bigr{)}^{1/(2d)}\geq\psi^{(0)}(X)e^{-\tau\Psi/2d}=4\gamma\psi^{(0)}(X)$ (25)
where the first bound comes from the inequality between arithmetic and geometric means, and the second bound comes from (24).
Partition $\mathcal{X}$ into sets of runs $\mathcal{X}^{\text{Y}},\mathcal{X}^{\text{N}}$ which output YES and NO respectively. By hypothesis of the lemma, we have ${\mathbb{P}}^{(r)}(\mathcal{X}^{\text{Y}})<\gamma$ for all $r\in R$. Using Eq. (25), this implies
${\mathbb{P}}^{(0)}(\mathcal{X}^{\text{Y}})=\negthinspace\int_{{\cal X}^{\text{Y}}}\negthinspace\psi^{(0)}(X)\>\operatorname{d}\!{\mu}^{\mathfrak{A}}X\leq\frac{1}{8d\gamma}\sum_{r\in R}\int_{{\cal X}^{\text{Y}}}\negthinspace\psi^{(r)}(X)\>\operatorname{d}\!{\mu}^{\mathfrak{A}}X=\frac{1}{8d\gamma}\sum_{r\in R}{\mathbb{P}}^{(r)}(\mathcal{X}^{\text{Y}})\leq\frac{1}{8d\gamma}\sum_{r\in R}\gamma=\frac{1}{4}.$
By hypothesis, ${\mathbb{P}}^{(0)}(\mathcal{X}^{\text{N}})<\gamma$. So $\mathfrak{A}$ runs for more than $\tau$ timesteps on $c^{(0)}$ with probability at least $1-{\mathbb{P}}^{(0)}(\mathcal{X}^{\text{Y}})-{\mathbb{P}}^{(0)}(\mathcal{X}^{\text{N}})\geq 1-\gamma-\frac{1}{4}$; by our assumption that $\gamma\in(0,\tfrac{1}{2})$, this is at least $\tfrac{1}{4}$. The expected cost is thus at least $\frac{1}{4}\cdot\tau=\Omega(\frac{d\log(1/\gamma)}{\Psi})$. ∎
@@
If $z(\beta_{\min},\beta_{\max})\leq q-2\kappa$, then $z^{(r)}(\beta_{\min},\beta_{\max})\leq q$ for all $r=\pm 1,\dots,\pm d$. Moreover, in this case:
(a) If $|z(\beta_{\min},\beta_{\max})-z^{(r)}(\beta_{\min},\beta_{\max})|>2\varepsilon$ for all $r=\pm 1,\dots,\pm d$, then algorithm for ${P^{{\varepsilon}}_{\tt ratio-point}}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})$ on $c^{(0)}$.
(b) If each $r\in\{1,\dots,d\}$ has values $x,y\in\mathcal{F}$ with $\Delta(x),\Delta(y)\geq\delta e^{2\kappa}$ and $|\lambda^{(r)}_{x}-\lambda^{(r)}_{y}|>2\varepsilon$, then any algorithm for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})$ on $c^{(0)}$. (We call $x,y$ the witnesses for instances $c^{(\pm r)}$).
@@
As $|\log c^{(r)}_{x}-\log c_{x}|\leq\kappa$ for all $x$, we have $\frac{Z^{(r)}(\beta)}{Z(\beta)}\in[e^{-\kappa},e^{\kappa}]$ and $\mu^{(r)}_{\beta}(x)=\frac{c_{x}^{(r)}e^{\beta x}}{Z^{(r)}(\beta)}\geq\mu_{\beta}(x)e^{-2\kappa}$. So $z^{(r)}(\beta_{\min},\beta_{\max})\leq q$ for all $r=\pm 1,\dots,\pm d$. In particular, the upper-bound value $q$ is valid for all the instances $c^{(r)}$.
In case (a), an algorithm for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ can distinguish $c^{(0)}$ from $c^{(\pm 1)},\dots,c^{(\pm d)}$: namely, given a solution $\hat{Q}(\beta_{\max})$ for ${P^{{\varepsilon}}_{\tt ratio-point}}$, we output YES if and only if $|\log\hat{Q}(\beta_{\max})-\log Q^{(0)}(\beta_{\max})|\leq\varepsilon$.
In case (b), an algorithm for ${P^{{\varepsilon}}_{\tt ratio-point}}$ can distinguish $c^{(0)}$ from $c^{(\pm 1)},\dots,c^{(\pm d)}$: namely, given a solution $\hat{c}$ to ${P^{\delta,\varepsilon}_{\tt count-weak}}$, we output YES if and only if for all $r=1,\dots,d$ the witnesses $x,y$ satisfy $|\log\frac{\hat{c}_{x}}{\hat{c}_{y}}-\log\frac{c_{x}}{c_{y}}|\leq\varepsilon$. Note that $\Delta^{(\pm r)}(x),\Delta^{(\pm r)}(y)\geq\delta$. ∎
By applying Corollary 46 to carefully constructed instances, we will show the following results which, in turn, immediately imply Theorems 2 and 4:
@@
Let $n>n_{0},q>q_{0},\varepsilon<\varepsilon_{0},\delta<\delta_{0}$ for certain absolute constants $n_{0},q_{0},\varepsilon_{0},\delta_{0}$. There are problem instances $\mu$ which satisfy the given bounds $n$ and $q$ such that:
@@
(c) ${P^{{\varepsilon}}_{\tt ratio-point}}$ requires cost $\Omega(\frac{\min\{q,n^{2}\}}{\varepsilon^{2}}\log\frac{1}{\gamma})$, and $\mu$ is log-concave.
@@
(e) ${P^{{\varepsilon}}_{\tt ratio-point}}$ requires cost $\Omega(\frac{q}{\varepsilon^{2}}\log\frac{1}{\gamma})$.
8.1 Bounds in the log-concave setting
To begin, observe that the lower bound of $\Omega(\frac{1}{\delta\varepsilon^{2}}\log\tfrac{1}{\gamma})$ for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ is rather trivial, and does not really depend on properties of Gibbs distributions. We summarize it in the following result:
@@ 8.1 Bounds in the log-concave setting
So it suffices to consider $\delta=\Omega(1)$, and we turn to the bounds in terms of $n$ and $q$. For this construction, take $d=1,\beta_{\min}=0,\beta_{\max}=n$, set $c_{0},\dots,c_{n}$ to be the coefficients of the polynomial $g(x)=\prod_{k=0}^{n-1}(e^{k}+x)$, and set $\lambda^{(1)}_{k}=10k\varepsilon/n$ for $k=0,\dots,n$.
Since $g(x)$ is a real-rooted polynomial, and counts $c^{(\pm 1)}$ are derived by an exponential shift, the counts $c^{(0)},c^{(+1)},c^{(-1)}$ are log-concave [5]. Due to the polynomial representation of the counts as $\prod_{k=0}^{n-1}(e^{k}+x)=c_{0}+c_{1}x+\dots+c_{n}x^{n}$, we have an elegant formula for the partition ratio function:
$Z(\beta)$ $=\sum_{x}c_{x}e^{\beta x}=g(e^{\beta x})=\prod_{k=0}^{n-1}(e^{k}+e^{\beta})$
$Z^{(\pm 1)}(\beta)$ $=\sum_{x}c_{x}e^{\pm\nu}e^{\beta x}=Z(\beta\pm\nu)\qquad\text{ for $\nu=10\varepsilon/n$}.$
For intuition, consider independent random variables $X_{0},\dots,X_{n-1}$, wherein each $X_{i}$ is Bernoulli-$p_{i}$ for $p_{i}=\frac{e^{\beta}}{e^{i}+e^{\beta}}$. Then $\mu_{\beta}$ is the probability distribution on the sum $X=X_{0}+\dots+X_{n-1}$, and distributions $\mu^{(\pm 1)}$ are derived by slightly shifting the probabilities $p_{i}$.
@@
$\mu_{\beta_{\min}}(0)$ $=\frac{c_{0}}{Z(0)}=\prod_{k=0}^{n-1}\frac{e^{k}}{e^{k}+1}\geq\prod_{k=0}^{n-1}\exp(-e^{-k})\geq\exp(-\sum_{k=0}^{\infty}e^{-k})=e^{-e/(e-1)}\geq 1/5.$
@@
For $n>n_{0},q>q_{0},\varepsilon<1/100$, where $n_{0},q_{0}$ are absolute constants, the instance can be constructed so that both ${P^{{\varepsilon}}_{\tt ratio-point}}$ or ${P^{1/10,\varepsilon}_{\tt count-weak}}$ require cost $\Omega(\frac{\min\{q,n^{2}\}\log\frac{1}{\gamma}}{\varepsilon^{2}})$.
@@
Here $\kappa=n\nu=10\varepsilon$. By reducing $n$ as needed, and taking $n_{0},q_{0}$ sufficiently large, we may suppose that $10100<n<\sqrt{q}/$ 10. In this case, we use Corollary 46 for both problems; with $z(\beta_{\min},\beta_{\max})+2\kappa\leq n^{2}+1/5\leq q$ for $q_{0}$ sufficiently large.
For ${P^{{\varepsilon}}_{\tt ratio-point}}$ we have $|z(\beta_{\min},\beta_{\max})-z^{(\pm 1)}(\beta_{\min},\beta_{\max})|>2\varepsilon$ as required. For ${P^{1/10,\varepsilon}_{\tt count-weak}}$ take witnesses $x=0,y=n$; we have $\Delta(x),\Delta(y)\geq 1/5\geq 1/10\cdot e^{2\kappa}$ and $|\lambda^{(1)}_{x}-\lambda^{(1)}_{y}|=n\nu=10\varepsilon$ as required. So both problems require cost $\Omega(\frac{\log(1/\gamma)}{\Psi})=\Omega(\frac{\log(1/\gamma)n^{2}}{\varepsilon^{2}})$. ∎